DVITODVI(1) DVITODVI(1) NAME dvitodvi−rearrangepagesinaDVIfile SYNOPSIS dvitodvi[−q][−iinfile][−ooutfile][−wwidth][−hheight]pagespecs[infile[outfile]] DESCRIPTION DvitodvirearrangespagesfromaDVIfile,creatinganewDVIfile.Dvitodvicanbeusedtoperformalargenumberofarbitraryre-arrangementsofDocuments,includingarrangingforprinting 2-up,4-up,booklets,reversing,selectingfrontorbacksidesofdocuments,scaling,etc. pagespecsfollowthesyntax: pagespecs=[modulo:][mag@]specs specs=spec[+specs][,specs] spec=[-]pageno[(xoff,yoff)] moduloisthenumberofpagesineachblock.Thevalueofmoduloshouldbegreaterthan0;the defaultvalueis1.Theoptionalmagparameteradjuststhefile’smagnificationbythemagnificationspecified.Themagnificationandmodulocanbespecifiedinreverseorderifdesired.specsare thepagespecificationsforthepagesineachblock.Thevalueofthepagenoineachspecshouldbe between0(forthefirstpageintheblock)andmodulo-1(forthelastpageineachblock)inclusive. Theoptionaldimensionsxoffandyoffshiftthepagebythespecified(positive)amount.xoffand yoffareinPostScript’spoints,butmaybefollowedbytheunitscmorintoconverttocentimetresorinches,ortheflagworhtospecifyasamultipleofthewidthorheight.Iftheoptional minussignisspecified,thepageisrelativetotheendofthedocument,insteadofthestart. Ifpagespecsareseparatedby+thepageswillbemergedintoonepage;iftheyareseparatedby theywillbeonseparatepages.Ifthereisonlyonepagespecification,withpagenozero,the pagenomaybeomitted. The−woptiongivesthewidthwhichisusedbythewdimensionspecifier,andthe−hoption givestheheightwhichisusedbythehdimensionspecifier.Thesedimensionsarealsoused(after scaling)tosettheclippingpathforeachpage. Dvitodvinormallyprintsthepagenumbersofthepagesre-arranged;the−qoptionsuppresses this. EXAMPLES Thissectioncontainssomesamplere-arrangements.Toputtwopagesononesheet(oflandscaped A3paper),thepagespectouseis: 2:0+1(21cm,0) ToreduceanA3pagetoA4size,thepagespectouseis: 700@(-.3in,-.3in) Toselectalloftheoddpagesinreverseorder,use: 2:-0 AUTHOR AngusDuggan,fromdviselect(1),byChrisTorek,UniversityofMaryland SEEALSO dviconcat(1),dvibook(1),dviselect(1),latex(1),tex(1) MC-TeXUser’sGuide 1 DVITODVI(1) DVITODVI(1) TheTeXbookDvitodvidoesnotadjustsomeoftheparametersinthepostamble.Themaybea problemifthesevaluesareusedtosizecertainstructuresintheoutputconversionprograms.